### v1

# Exploring Knowledge Tracing in Tutor-Student Dialogues

Abstract
Recent advances in large language models (LLMs) have led to the development of artificial intelligence (AI)-powered tutoring chatbots, showing promise in providing broad access to high-quality personalized education. Existing works have primarily studied how to make LLMs follow tutoring principles but not how to model student behavior in dialogues. However, analyzing student dialogue turns can serve as a formative assessment, since open-ended student discourse may indicate their knowledge levels and reveal specific misconceptions. In this work, we present a first attempt at performing knowledge tracing (KT) in tutor-student dialogues. We propose LLM prompting methods to identify the knowledge components/skills involved in each dialogue turn and diagnose whether the student responds correctly to the tutor, and verify the LLM’s effectiveness via an expert human evaluation. We then apply a range of KT methods on the resulting labeled data to track student knowledge levels over an entire dialogue. We conduct experiments on two tutoring dialogue datasets, and show that a novel yet simple LLM-based method, LLMKT, significantly outperforms existing KT methods in predicting student response correctness in dialogues. We perform extensive qualitative analyses to highlight the challenges in dialogue KT and outline multiple avenues for future work.
Keywords Knowledge Components  $\cdot$ Knowledge Tracing  $\cdot$ Large Language Models  $\cdot$ Tutoring dialogues
1 Introduction
Tutoring, which mostly takes the form of dialogues between tutors and students, has proven to be highly effective at improving student learning outcomes [36]. Recent advances in generative artificial intelligence (AI), especially large language models (LLMs), have led to the development of AI-powered tutoring chatbots, such as Khan Academy’s Khanmigo [22] and Carnegie Learning’s LiveHint [4]. By wrapping LLMs around learning content such as textbooks and practice questions, these automated chatbots can interact with students in real time, potentially reaching a larger number of students and alleviating the burden on human tutors.
To use LLMs as tutors, it is important to i) ensure they behave like human tutors and ii) control the safety of their output. Since most of existing LLM-powered tutoring chatbots are developed by either fine-tuning on tutoring dialogue transcripts or prompt engineering, plenty of works over the last few years have studied how to make LLMs behave according to effective tutoring principles. The work in [53] studies the latent decision-making process behind human tutor utterances and shows that adopting decisions by human experts can benefit LLMs in tutoring. The work in [33] identifies tutor moves when they interact with an LLM-powered simulated student agent and shows that move annotation data can make LLMs become better tutors. The work in [17] generates large amounts of synthetic tutoring data via AI roleplay and shows that fine-tuning on this data, in addition to data collected from human tutors and students, makes LLMs more effective in pedagogy. The work in [14] proposes a way to measure tutor uptake, i.e., acknowledging what students have said, which is another beneficial strategy for LLMs to adopt. The work in [42] attempted to summarize tutor utterances into actionable feedback for students using LLMs, but found that the resulting messages were unable to outperform those written by real human teachers.
One important observation is that existing works have almost exclusively studied the tutor turns in dialogues rather than the student turns. A possible reason is that student turns can be noisy in nature, partially due to students not fully mastering the required knowledge, resulting in errors that are harder to analyze than tutor utterance. Take Google’s LearnLM [17] for example: student dialogue turns in their AI roleplay-based synthetic training data are generated by prompting an LLM to make an error that a real student would make. This error information is then fed to the tutor LLM to help the tutor model identify the error and provide corresponding feedback. However, there are two key limitations to this approach: First, assuming LLMs can behave like real students may not be reliable. As recent work [15] shows, LLMs are ineffective at anticipating, parsing, and following flawed reasoning. Second and more importantly, real open-ended student discourse is highly valuable since it may indicate flaws in student knowledge and even reveal specific misconceptions. Therefore, analyzing student dialogue turns can be a form of formative assessment to help tutors and AI-powered chatbots understand student progress and provide more targeted feedback. It can also potentially improve simulated student agents [21] to make them better at reflecting real student behavior, which can, in turn, improve the training of tutor chatbots through reinforcement learning [18, 43].
1.1 Contributions
In this work, we propose dialogue knowledge tracing, a novel task, which seeks to analyze student discourse using the knowledge tracing (KT) framework [9]. We enable KT on dialogues via a two-step process: i) we use pre-trained LLMs, specifically GPT-4o [37], to annotate existing dialogues with student correctness and knowledge component (KC) tags using Common Core standards, and ii) we train KT models on the annotated data to estimate student knowledge at the dialogue-turn level. We introduce two novel KT methods that leverage the textual content in dialogues: an LLM-based KT method (LLMKT), which fine-tunes the open-source Llama 3 model [34] for the KT objective, and DKT-Sem, which replaces the embeddings of DKT [41] with semantic text embeddings. Using two existing tutor-student math dialogue datasets, we show that i) our new methods significantly outperform existing KT methods, especially when little training data is available, and ii) all KT methods can accurately predict student correctness given enough training data. We also use qualitative analysis and learning curve visualizations to show that LLMKT learns meaningful knowledge state estimates. Finally, we show that GPT-4o’s dialogue annotations are accurate according to human experts.
We acknowledge many limitations of our work upfront since it is, to the best of our knowledge, the first at attempting to perform KT in tutoring dialogues. First, a “time step” in our KT setup corresponds to a single student turn in the dialogue, which may make knowledge states volatile; this setting differs from the scale of a time step in existing KT methods, which usually corresponds to a student responding to an assessment/practice question. We cannot explore setting an entire dialogue as a time step since no existing open-source tutoring dialogue dataset links students across dialogues. Second, our work is restricted to math since analyzing the KCs in math problems is challenging yet well-studied; it remains to be seen whether our approach of Common Core standard tagging generalizes to other subjects. Third, due to the smaller scale of tutoring dialogue datasets compared to those typically used in standard KT works, our methods need to be tested on larger-scale data. One dataset we use in our experiments contains partial dialogues that are cut short for LLM evaluation purposes [35], which further adds to this issue. Therefore, we will publicly release our code [^1] https://github.com/umass-ml4ed/dialogue-kt so that researchers who own a large amount of dialogue data can experiment on their proprietary data.
2 Related Work
2.1 Knowledge Tracing
KT [9] is a well-studied task in the student modeling literature. It studies the problem of breaking down student learning into a series of time steps practicing certain KCs and using the correctness of each response step to track student knowledge. Classic Bayesian knowledge tracing methods [39] use a latent binary-valued variable to represent whether a student masters a KC or not. Factor analysis-based methods [6, 40] use features in addition to latent ability parameters in a logistic regression setting to predict student response correctness. Since neural networks became popular, plenty of deep learning-based KT methods were developed. These models have limited interpretability since student knowledge is represented as hidden states in complex, deep neural networks. Most of these methods use long short-term memory networks [19] or variants [41, 48], with other variants coupling them with memory augmentation [58], graph neural networks [56], attention networks [16, 38], or pre-trained word embeddings to leverage question text information [28].
Since LLMs started to find success in natural language applications, plenty of researchers have attempted to use them in some ways for KT. Some use the neural network architecture of LLMs and train on student response data [57]. Others use LLMs to represent questions and responses in certain domains such as language learning [11, 50] and computer science education [24, 27]. More recently, [26] attempted prompting-based KT that directly asks LLMs to represent questions and student knowledge states as text. Despite some success in domains such as math, it remains to be seen whether these models for question response data can be useful in text-heavy KT situations such as dialogues.
2.2 Knowledge Component Assignment and Discovery
There exist a lot of prior works on automated KC tagging and discovery, including the notable Q-matrix method [3], which assigns KCs to questions using student response data. More recently, methods such as [47] also learn latent KCs by training deep learning models on KT data and enforcing priors from pedagogical theory. While these methods effectively learn new KCs, they do not produce textual descriptions of KCs, making them difficult to interpret. Several recent works have used LLMs to assign expert-written KCs to math problems via prompting [31, 55], fine-tuning [46], and reinforcement learning [25]. However, none have attempted assigning KCs in dialogues. A prior work has attempted to tag KCs in dialogue turns [10], although using a manual approach that cannot be scaled.
2.3 Dialogue Analysis
As summarized above, existing works on dialogue analysis have primarily focused on analyzing tutor behavior. The works in [13, 51] automatically classify the “talk moves” that tutors and students make via fine-tuned LLMs, the work in [14] automatically measures tutors’ “uptake” behavior in dialogues, and the work in [1] uses talk moves in dialogues to predict student performance in post-tests. To the best of our knowledge, few prior works have systematically studied the problems of KC tagging, student behavior analysis, and particularly KT, in dialogues.
3 KT in Dialogues
We now formulate the dialogue KT framework. Our goal is to develop methods that can i) estimate student knowledge on KCs involved in each dialogue turn and ii) use these estimates to predict the correctness of student dialogue turns. Intuitively, we can view tutor-student dialogues as sequences of (formative) assessment questions. Each tutor turn acts as a question related to some learning objectives and one or more KCs. These turns can correspond to many different tutor moves: asking a clarification question, identifying a student error, providing actionable feedback, prompting the student to think in a Socratic way, or offering encouragement [14, 52]. Each student turn, on the other hand, is a correct or incorrect response to the task posed by the tutor, which reflects their thinking process and offers insights into their knowledge state. A pair of tutor-student dialogue turns corresponds to a time step in KT. Therefore, we split the dialogue KT framework into three stages: i) identify the KCs involved in each student dialogue turn, ii) classify whether the student response is correct or incorrect, and iii) apply new and existing KT methods on the resulting data.
Formally, we define each dialogue $d$ to be a sequence of alternating tutor turns, $t$, and student turns, $s$, i.e., $d=(s_{0},t_{1},s_{1},\ldots,t_{M},s_{M})$, where $M$ is the total number of turn pairs for that dialogue and $s_{0}$ is present if the student initiates the dialogue. Each relevant turn pair is assigned a correctness label $y_{j}\in\{0,1\}$, where $j\in\{1,\ldots,M\}$ is the turn pair index, $y_{j}=1$ means that $s_{j}$ is a correct response to $t_{j}$, and $y_{j}=0$ means that it is incorrect. In practice, we also use $y_{j}=\text{na}$ in cases where correctness is not well-defined for $t_{j}$, e.g., the turn is off-topic or the tutor is offering emotional support. Each tutor turn is also associated with a set of KCs, $\mathcal{C}_{j}=\{c_{j1},\ldots,c_{jK}\}$, where $K$ is the number of KCs for that turn. These KCs are related to learning objectives outlined by the tutor, which lists knowledge elements that the student must posses to respond correctly. In this work, we use the Common Core math standards [8] as the KCs since they are i) standard and widely adopted and ii) textual in nature and are interpretable to both humans and LLMs. For irrelevant dialogue turns that we do not consider, $\mathcal{C}_{j}=\emptyset$.
In a dialogue turn, we use $z_{jk}\in\{0,1\}$ to denote the student’s unobserved, binary-valued latent mastery status of the KC, $c_{jk}$, that is involved in the current step. $z_{jk}=1$ means that the student masters KC $k$ in dialogue turn $j$. Intuitively, the student needs to master all KCs involved in a turn to respond correctly to the task posed by the tutor. Formally, the probability of a correct student response at turn $j$ can be written as $P(y_{j}=1|z_{j1},\ldots,z_{jK})$. The latent KC mastery depends on the entire dialogue history, i.e., $P(z_{jk}=1)=\phi(c_{jk},t_{\leq j},s_{<j},y_{<j},\mathcal{C}_{<j})$, where $\phi$ returns the probability of the student mastering KC $c_{jk}$, and subscripts $<j$ and $\leq j$ correspond to all past student and (including the current) tutor dialogue turns, respectively. In the next section, we detail how we parameterize and train this function in practice.
4 Methodology
We now detail our two primary technical contributions. First, we detail how to prompt pre-trained LLMs to automatically annotate dialogue turns to collect the KC and response correctness labels needed for dialogue knowledge tracing. Second, we propose new KT methods for the dialogue KT task that incorporate the rich textual information in tutor-student dialogues. We show a demonstration of the dialogue KT framework and the LLMKT method in Figure 1.
Figure 1: Overview of the dialogue KT framework and LLMKT. For annotation, we provide a full dialogue to GPT-4o, which generates correctness labels and retrieves Common Core standards as KCs for each turn. For KT, LLMKT takes a dialogue up to the target turn and its KCs to predict mastery on each KC and response correctness.
4.1 Automated Dialogue Annotation
In order to perform knowledge tracing on dialogues, each student conversation turn must first be annotated with two labels: i) the correctness of the student response, and ii) a list of KCs involved in the turn. While it is possible for human experts to provide these annotations, doing so would be incredibly time consuming and difficult to scale. Following recent works that have had success using LLMs to annotate data in educational tasks [45, 53], we use a state-of-the-art LLM, GPT-4o [37], to perform the annotation. We conduct the two annotation tasks separately since we find that doing so leads to improved accuracy compared to combining them. We use a simple, zero-shot chain-of-thought prompting approach and leave exploration of more advanced prompting approaches, such as few-shot prompting or self-reflection, for future work. To facilitate replication, we provide our prompts in an online Supplementary Material [^2] https://osf.io/873ms?view_only=f32472a6560649e692a4e2ba603ef52c. In Section 7, we conduct a human evaluation with math teachers to show that the GPT-provided annotations are mostly accurate and aligned with human judgement.
For correctness annotation, we provide the full dialogue as input to the LLM and instruct it to first summarize each turn in the dialogue in a CoT way, before labeling the correctness of each turn. We specifically instruct the LLM to “summarize” to leverage the fact it is pre-trained extensively on summarization tasks; this approach is inspired by recent work [32] that has shown that such an approach is effective for tasks such as code editing.
For KC annotation, we adapt the recursive tagging algorithm from [31], which tags math word problems with Common Core math standards. The algorithm uses hierarchically organized standards from the Achieve the Core [^3] https://achievethecore.org/ (ATC) coherence map, where standards are grouped into mid-level “clusters”, which are further grouped into high-level “domains”. To assign standards to a dialogue, we first provide the LLM with the full dialogue and the list of all 11 domains, and ask it to summarize the dialogue before selecting all domains that are relevant to the learning objectives. We then take the union of all clusters that are children of the selected domains and prompt the LLM again, this time asking it to summarize at the turn level but still select relevant clusters at the dialogue level. Finally, we take the union of all standards that are children of the selected clusters and prompt the LLM a final time, having it summarize each turn and assign standards at the turn level. We treat each Common Core math standard as a KC.
We note that by annotating a full dialogue with a single prompt, knowledge tracing in real-time becomes impossible since KCs for turn $j$ are assigned with $s_{j}$ and future information in context. This limitation could be alleviated in practice by annotating dialogues one turn at a time without seeing $s_{j}$ when labeling $C_{j}$. However, since doing so would significantly increase the number of API calls and thus the cost of annotation, and we are not currently studying the real-time setting, we leave exploration of individual turn tagging and real-time applications for future work.
4.2 LLMKT
Another feature that makes dialogue KT different than the regular KT task is that most dialogue turns (at least 58%) are associated with more than one KC; see Table 1 for details. In contrast, most existing KT methods are developed for questions that cover a single KC; solutions that can deal with the occasional multi-KC question in existing KT datasets include i) converting a single question-response pair with multiple KCs to multiple pairs for different “questions”, each covering a single KC, and ii) treating a unique combination of multiple KCs as a new KC. However, due to the prevalence of dialogue turns tagged with multiple KCs, we found that these solutions are not applicable to the KT-in-dialogues setting. Therefore, we resort to an “AND”-type model, where the student is likely to respond to a tutor-posed task correctly only if they master all KCs involved in the turn. Formally, we model the probability of a correct student response at a turn as the product of the KC mastery across ones involved in this turn, i.e.,
$P(y_{j}=1)=\textstyle\prod_{k=1}^{K}P(z_{jk}=1).$ (1)
We now detail our proposed LLM-based KT method, which we refer to as LLMKT. One can, in theory, estimate student knowledge states $z_{jk}$ by simply providing the text of the dialogue up to $t_{j}$ to a pre-trained LLM, and prompting it to predict whether the student masters the KCs $c_{jk}$ for each $k\in[1,K]$. However, since LLMs are not pre-trained on the specific KT task, we cannot expect them to do well in this “zero-shot” setting. Therefore, we adopt a fine-tuning approach using an open-source, decoder-only LLM, such as Llama 3 [34], for the KT task. We model the estimated probability of a student mastery $\hat{z}_{jk}$ as
$\hat{z}_{jk}=P_{\theta}(z_{jk}=1)=\frac{\exp(v^{\mathcal{T}}_{\theta}(t_{\leq j},s_{<j},c_{jk}))}{\exp(v^{\mathcal{T}}_{\theta}(t_{\leq j},s_{<j},c_{jk}))+\exp(v^{\mathcal{F}}_{\theta}(t_{\leq j},s_{<j},c_{jk}))},$ (2)
where $\theta$ represents the model parameters, and $v^{\mathcal{T}}_{\theta}(\cdot)$ and $v^{\mathcal{F}}_{\theta}(\cdot)$ return the output logits of the LLM at the “True” and “False” vocabulary tokens, respectively. In other words, with i) the entire conversation history and ii) a textual prompt instructing the LLM to estimate the student’s current KC mastery level as input, we use an LLM to parameterize this estimate through two token generation probabilities. In practice, to improve model efficiency, we pack all KCs in $C_{j}$ into a single prompt; we use customized attention masks and position embeddings to ensure that KCs do not attend to each other and their mastery levels are estimated independently. We do not explicitly provide KCs or correctness labels for previous turns due to memory constraints, but observe that LLMs can infer them from the context.
We fine-tune the LLM by maximizing the likelihood of all observed student response correctness labels, or equivalently, minimizing the binary cross entropy (BCE) loss between the ground truth correctness labels $y_{j}$ and the predicted student correctness probability $\hat{y}_{j}$ as
$\hat{y}_{j}=\textstyle\prod_{k=1}^{K}\hat{z}_{jk},\quad\mathcal{L}(d;\theta)=-\textstyle\sum_{j=1}^{M}y_{j}\log\hat{y}_{j}+(1-y_{j})\log(1-\hat{y}_{j}).$ (3)
While this method is relatively simple, it has many advantages. First, because we start with a powerful pre-trained LLM, LLMKT is able to leverage existing textual analysis capabilities that are not available in other KT architectures. Second, we do not add any new parameters to the pre-trained model, likely reducing the need for large amounts of training data. Third, since LLMKT operates on textual content rather than question/KC embeddings that need to be learned from scratch, it should be more generalizable to new KCs that are not seen during training.
4.3 DKT-Sem
In addition to LLMKT, we experiment with an alternative, simpler KT method for dialogues, which we can be viewed as a strong baseline. We slightly modify the deep KT (DKT) model [41] structure to use semantic embeddings of the textual content in dialogues, resulting in a method we term DKT-Sem. We simply use a learnable projection matrix to map text embeddings of every dialogue turn to the latent knowledge space in DKT, and similarly predict masteries with a learnable projection from the latent space to the text embeddings of KCs. We believe that this simple method can result in better generalization across KCs and dialogues compared to DKT, which learns question and KC representations from scratch. Formally, we model KC mastery probability as
$\mathbf{t}_{j}=\operatorname{S-BERT}(t_{j}),\quad\mathbf{s}_{j}=\operatorname{S-BERT}(s_{j}),\quad\mathbf{c}_{j}=\textstyle\frac{1}{K}\sum_{k=1}^{K}\operatorname{S-BERT}(c_{jk}),\quad\mathbf{y}_{j}=\operatorname{emb}({y_{j}})$ (4)
$\mathbf{x}_{j}=\tanh([\mathbf{t}_{j};\mathbf{s}_{j};\mathbf{c}_{j}]\mathbf{W})\oplus\mathbf{y}_{j},\quad\mathbf{h}_{j}=\operatorname{LSTM}(\mathbf{x}_{\leq j}),\quad\hat{\mathbf{z}}_{j+1}=\sigma(\mathbf{h}_{j}\mathbf{B}\mathbf{C}^{T}).$ (5)
The input at each time step is created by first computing the sentence BERT (S-BERT) [44] embeddings of the teacher turn, $\mathbf{t}_{j}$, student turn, $\mathbf{s}_{j}$, and element-wise average embeddings of the KCs, $\mathbf{c}_{j}$, respectively. Following DKT, response correctness is mapped to a learnable embedding, $\mathbf{y}_{j}$, which is then added element-wise to a linearly transformed version (through $\mathbf{W}$) of the concatenated tutor turn/student turn/KC embeddings, resulting in the input vector $\mathbf{x}_{j}$. Finally, the vector of estimated KC masteries at the next turn, $\hat{\mathbf{z}}_{j+1}$, are computed via a bilinear projection, $\mathbf{B}$, between the LSTM hidden state, $\mathbf{h}_{j}$, and the matrix of all KC S-BERT embeddings, $\mathbf{C}$, followed by the sigmoid function. We train DKT-Sem using the same objective as LLMKT (Eq. 3). Due to the smaller model size, DKT-Sem is easier to train than LLMKT, especially when access to computational resources to support LLM fine-tuning is limited.
5 Experiments
5.1 Datasets
We experiment with two datasets, CoMTA and MathDial, to study the dialogue KT task and validate the proposed KT methods. We acknowledge that there are many other datasets available, such as the ones in [13, 52, 53] and ones not focusing on math education. We select CoMTA since it is recent and highly reflective of AI-student dialogues that are becoming prevalent, and we select MathDial since it is relatively large-scale. However, our framework should be applicable to other math tutoring dialogue datasets as well. We show statistics for these datasets in Table 1.
Table 1: Statistics for the CoMTA and MathDial datasets, from left to right: total number of dialogues, total number of turn pairs with correctness labels, percentage of turn pairs where the student response is correct, total number of unique KCs in the dataset, average number of unique KCs per dialogue, average number of KCs per turn pair, and percentage of turn pairs with more than one KC.
Dataset Num. Dia. Num. Labels Correct Num. KCs Avg. KCs/Dia. Avg. KCs/Turn Turns >1 KC
CoMTA 153 623 57.78% 164 3.20 1.94 58.75%
MathDial 2,823 13,200 49.86% 145 4.16 2.32 82.89%
The CoMTA dataset [35] contains 188 dialogues between human students and Khanmigo, Khan Academy’s GPT-4-powered tutor. In most dialogues, students ask for help with a specific topic or problem(s) that they do not know how to solve, and work through one (or more) problems with the tutor throughout the dialogue. The tutor is primed with a series of customized prompts developed by Khan Academy to behave like real human tutors: to use the Socratic method, frequently asking questions at almost every dialogue turn, and to not directly reveal correct answers to students. Each dialogue centers around a math subject among Elementary, Algebra, Trigonometry, Geometry, and Calculus. However, since Common Core does not contain Calculus standards, we remove Calculus dialogues, resulting in 153 remaining dialogues. Since no train/test split is defined, we perform a 5-fold cross-validation in experiments.
The MathDial dataset [33] contains 2,848 dialogues between students simulated by GPT-3.5 and crowd workers role-playing as human tutors. Each dialogue is in the context of an incorrect “student” solution to a math problem, and the tutor’s goal is to help the student arrive at the correct solution by helping them identify and correct their errors. We use the original train/test split, containing 2,235/588 dialogues, respectively, after removing a total of 25 dialogues, where GPT-4o’s annotation of student response correctness and turn KCs failed.
5.2 KT Methods
In addition to LLMKT and DKT-Sem, we also evaluate the performance of a suite of widely-used KT methods on the dialogue KT task. Our goal is to test whether existing KT methods, developed for a slightly different use case, are applicable to our setting and how they compare to methods designed for dialogue KT.
We evaluate multiple popular deep learning-based KT methods: DKT [41], DKVMN [58], AKT [16], SAINT [7], and simpleKT [29], all sourced from the pyKT code repository [30]. There are two key differences between these existing KT methods and our proposed KT methods for dialogues: i) they learn new embeddings for questions, KCs, and response correctness without leveraging any textual information, and ii) they are designed to analyze questions that cover a single KC. Since most of the dialogue turns contain at least two KCs in the CoMTA and MathDial datasets, we convert each turn into $K$ “pseudo-turns” that are each associated with one of the KCs, $c_{jk}$; all such pseudo-turns are associated with the same correctness label, $y_{j}$, of that turn. This modification follows common practice in the KT literature [16, 29]. We predict a single $\hat{\mathbf{z}}_{j}$ vector for each turn pair, using the output of the most recent time step that does not take $y_{j}$ as input, to avoid leaking the correctness label across pseudo-turns. We note that predicting a vector instead of a single probability requires altering the size of the output projection for all methods except DKT. We train all methods using the same objective as LLMKT and DKT-Sem (Eq. 3).
Bayesian KT (BKT) [9] is a classical KT method that uses a hidden Markov model to learn interpretable, per-KC learning acquisition parameters and use them to predict changing mastery levels for a student. Since BKT is designed to model data where each question is only assigned a single KC, we similarly expand turns into “pseudo-turns” as described above. Due to the rigidity of BKT, we change the setup slightly compared to other methods. At train time, we train on each pseudo-turn label directly rather than aggregating predictions at the true turn level. At test time, we collect predictions on all pseudo-turns and take their mean to get the predicted correctness for a real turn; taking a product, as the other methods do, would yield very low probabilities because BKT does not perform a product at train-time. In contrast to the other methods, this setup of predicting labels at the pseudo-turn level does not leak correctness labels because information does not interact across KCs in BKT. We use pyBKT [2] to train and generate predictions.
5.3 Metrics and Experimental Setup
To quantitatively evaluate the effectiveness of KT methods in the dialogue KT task, we employ 3 popular metrics: i) accuracy (Acc.), the portion of predicted labels that match the ground truth after rounding $\hat{y}_{j}$ to 0 or 1, ii) the area under the receiver operating characteristic curve (AUC), the most widely used metric in the KT literature, and iii) F1, the harmonic mean between binary precision and recall, again using rounded predictions. We compute these metrics on all correctness labels except for the first label in each dialogue, since existing KT methods are unable to make educated estimates without prior labels in context.
We implement LLMKT using the Llama-3.1-8B-Instruct model [34] in the Huggingface Transformers library [54] and adapt parameters in fine-tuning with LoRA [20]. We implement DKT-Sem using PyTorch and use the all-mpnet-base-v2 model to generate S-BERT embeddings. During training, we randomly take 20% of the dialogues in the train set to use as a validation set and perform early stopping by calculating the loss on this set. We then evaluate all metrics on the test set, and for CoMTA repeat training and testing for all 5 folds and report the average and standard deviation of each metric across folds. We train on a single NVIDIA RTX A6000 GPU.
Before computing final results, we perform a hyperparameter search to find optimal settings for each method. We perform an iterative search for LLMKT and a grid search for other methods, keeping the checkpoint with the highest AUC on the validation set. For LLMKT, we perform the search on the CoMTA dataset and use the same parameters on MathDial to reduce training time, and for other methods we use the best hyperparameter values found for each dataset. For LLMKT, we search over learning rate $\in\{5\cdot 10^{-5},1\cdot 10^{-4},2\cdot 10^{-4},3\cdot 10^{-4}\}$ and LoRA’s rank $r\in\{4,8,16,32\}$, set batch size to 64 using gradient accumulation, gradient norm clipping to 1.0, and train for 5 epochs. For other KT methods, we search over learning rate $\in\{1\cdot 10^{-4},2\cdot 10^{-4},5\cdot 10^{-4},1\cdot 10^{-3},2\cdot 10^{-3},5\cdot 10^{-3}\}$ and embedding size $\in\{8,16,32,64,128,256\}$, set batch size to 64, train for 100 epochs, and set all other model-specific hyperparameters to the defaults in pyKT. For all methods, we train with the AdamW optimizer with a weight decay of $1\cdot 10^{-2}$. We found the optimal learning rate and $r$ of LLMKT to be $2\cdot 10^{-4}$ and 16, respectively; the optimal hyperparameters for existing KT methods varied across methods and datasets, but tended to have smaller embedding sizes on CoMTA and larger ones on MathDial since the former is much smaller in scale compared to the latter.
6 Experimental Results
6.1 Quantitative Results
Table 2 shows the quantitative performance of all KT methods on both datasets. We see that LLMKT significantly outperforms all other methods on both datasets, confirming its effectiveness in the dialogue setting. This advantage over other KT methods is likely due to the use of a pre-trained LLM to capture the rich textual content in dialogues. As a result, the LLM can estimate student knowledge after fine-tuning on student turn response correctness labels. We also see that DKT-Sem performs as well or better than existing deep learning KT methods even though it largely uses the same model architecture as DKT. This result further emphasizes the benefit of utilizing the textual dialogue context in some way, even if the original model architecture of existing KT methods remains largely unchanged.
Table 2: Student dialogue turn correctness prediction performance for all KT methods on both datasets, with best numbers in bold and second-best underlined. LLMKT significantly outperforms all other methods on both datasets.
CoMTA MathDial
Method Acc. AUC F1 Acc. AUC F1
BKT $51.02_{\pm 11.16}$ $52.50_{\pm 10.61}$ $53.41_{\pm 13.53}$ 60.71 64.19 56.71
DKT $46.76_{\pm 5.59}$ $50.70_{\pm 6.33}$ $32.52_{\pm 9.48}$ 57.18 60.13 53.45
DKVMN $46.78_{\pm 4.76}$ $53.37_{\pm 7.84}$ $28.30_{\pm 3.70}$ 57.58 60.43 57.82
AKT $48.06_{\pm 3.48}$ $47.94_{\pm 6.50}$ $40.05_{\pm 8.73}$ 58.74 59.81 51.91
SAINT $48.71_{\pm 7.87}$ $47.00_{\pm 9.01}$ $32.30_{\pm 7.47}$ 54.96 58.01 51.57
simpleKT $48.53_{\pm 4.12}$ $47.64_{\pm 5.22}$ $42.54_{\pm 5.22}$ 57.38 60.50 53.97
DKT-Sem $\underline{58.59}_{\pm 8.64}$ $\underline{62.65}_{\pm 4.49}$ $\underline{60.93}_{\pm 14.22}$ 57.78 61.13 57.07
LLMKT $\textbf{60.36}_{\pm 5.95}$ $\textbf{65.15}_{\pm 6.62}$ $\textbf{65.54}_{\pm 6.96}$ 65.09 70.71 63.89
On CoMTA, we see that all existing KT methods almost completely fail, unable to outperform the simple baseline of predicting the majority class label (which corresponds to an Acc. of $57.83_{\pm 5.08}$, AUC of $50.0_{\pm 0.00}$, and F1 of $58.87_{\pm 29.70}$). This result is expected since the size of the CoMTA dataset is small and neural network-based KT methods are data-hungry. Another important factor is that there are 164 unique KCs across only 623 labeled dialogue turn pairs, which is not enough for these methods to properly learn the KC embeddings or for BKT to learn the per-KC parameters. In contrast, LLMKT needs only minimal training data to reach acceptable accuracy, due to the underlying LLM being highly effective at understanding previous student dialogue turns. It can also leverage KC text information to generalize across KCs, which is especially beneficial for ones that do not occur frequently.
On MathDial, which has 21x more data than CoMTA, we see that existing KT methods perform significantly better (for reference, the majority baseline only gets an Acc. of $52.64$ since class labels are balanced). This result is expected and shows that they can perform the dialogue KT task when given enough training data. BKT outperforms neural network-based KT methods since even the larger MathDial dataset has short sequence length (less than 5 turns on average). Therefore, recurrent or attention neural networks that can capture long context history may not outperform simple models in this setting. We also see that DKT-Sem does not outperform existing KT methods, which suggests that the way it uses textual content only helps when little training data is given. However, LLMKT still maintains a large advantage over all other KT methods. This result suggests that even on large-scale datasets, prior dialogue history can still significantly improve KT performance, even when labeled student turn correctness data is abundant. However, this rich contextual information needs to be leveraged effectively, likely by an LLM-based KT method.
Finally, we see that performance of all KT methods is relatively low on the dialogue KT task, with a maximum of around $71\%$ AUC for student dialogue turn correctness prediction. This observation is in stark contrast to results on the standard KT task, where state-of-the-art KT methods can reach $80\%$ AUC for question response correctness prediction. This result suggests that dialogue KT is a challenging task, due to various reasons: First, the dialogues are relatively short, with only 4-5 labeled turn pairs on average, giving very little context for KT methods to leverage. For example, dialogues in CoMTA are cut short on purpose, to serve the goal in [35] on evaluating how accurate LLMs are at math. Therefore, we call upon researchers who have access to dialogues with more turns to validate our findings on their (possibly proprietary) data using our publicly available implementation. Second, student behavior in dialogues is more unpredictable compared to traditional KT, since there can be many diverse types of student responses to tutor turns. We believe that these challenges make dialogue KT a task worth studying in future work; it is likely that new KT methods need to be developed to improve performance on this task.
6.2 Qualitative Analysis
We now detail findings from our qualitative analysis, where we seek to i) identify challenges in the dialogue KT task and ii) understand why LLMKT has an advantage over existing KT methods. Therefore, we examine LLMKT’s predicted KC mastery and correctness on dialogues that are in the test set from our experiments on CoMTA.
We identify three primary challenges in the data that are unique to the dialogue KT task that are difficult for KT methods. First, the open-ended nature of dialogues means that student behavior is much less structured than in traditional assessment or practice settings. For example, when the tutor poses a question to the student, the student often i) does not answer directly but instead asks a clarifying question, ii) skips ahead to the final solution in the case of step-by-step problem solving, or iii) asks the tutor to give them a different problem. These student actions are inherently difficult to anticipate ahead of time since they define behavior beyond what can be captured by the KCs. As a result, there are many turns in these dialogues that are “irrelevant” to the KT task. Future work may develop methods to classify student intent and automatically identify these turns.
Second, since dialogues can be short (around 4 turn pairs on average), there is little historical context to estimate student knowledge from. In many dialogues, students are proficient in one aspect of problem solving, responding to several questions correctly at first, but then struggle when the tutor moves on to a new KC. Without sufficiently long student history, it is difficult to estimate KC mastery levels for ones that are seen for the first time. This problem is partially due to existing datasets not linking students over multiple dialogues. Future work may alleviate this problem by tracking students over multiple dialogues or including additional student-level information, such as pre-test scores.
Third, the varying number of KCs in different dialogue turns can lead to numerous false negatives and false positives, due to the product in our “AND”-type correct response likelihood model (Eq. 1). Even if the predicted KC mastery levels are high, the predicted correctness probability will be low for a turn will multiple KCs. On the contrary, turns with fewer KCs are more likely to be predicted as correct. Future work can explore other more sophisticated models for turns with multiple KCs, such as DINA [12] or Boolean logic analysis [23], to identify which KCs are more important to the correctness of a response rather than treating them equally.
We find that, unsurprisingly, LLMKT’s advantage over existing KT methods comes from leveraging nuanced details in the dialogue, rather than solely relying on previous correctness and KC labels. For the most part, the predicted mastery level correlates with previous turn correctness; as a result, dialogues with many correct or incorrect turns in a row are easy to predict for all KT methods. LLMKT excels in cases when it relies on textual information, rather than correctness labels, to adjust its KC mastery estimates. Table 3 shows one such example. At the first turn, the student responds incorrectly because they do not provide the answer the tutor was looking for. However, the response still shows some understanding of the topic; therefore, LLMKT increases, rather than decreases, the KC mastery prediction, despite the incorrect label for this turn. As a result, it correctly predicts that the student responds correctly to later turns while other KT methods cannot.
We also found that LLMKT tends to predict low mastery when a KC is seen for the first time in a dialogue, possibly through observing that students do not react well to the tutor shifting to new KCs mid-way through the dialogue. This observation also suggests that LLMKT can effectively use dialogue context to detect when new KCs are introduced. Overall, these observations suggest that there are numerous advantages in exploiting the textual content in dialogues for the dialogue KT task, rather than just relying the correctness and KC labels; LLMKT seems to be an encouraging step in this direction for future KT methods.
Table 3: An example dialogue with student turn correctness labels, KC labels, and LLMKT’s predicted mastery level for each KC. Even though the student’s response at the first turn is not fully correct, it suggests that they do have some knowledge on underlying topic, causing LLMKT to increase its mastery predictions in the next steps as the student responds correctly.
Speaker Turn Text Knowledge Components - $\hat{z}_{jk}$
Student how do I multiply this? 7x $\times$ 8
Tutor Ah, you’re asking about multiplication! Let’s break it down. What do you get when you multiply $7x$ by $8$? 1) Write, read, and evaluate expressions in which letters stand for numbers. $0.6792$
Student is it possible to multiply the 8 and 7?
Tutor You’re thinking correctly! When we perform multiplication, we can rearrange the numbers in any order. So, we can first multiply 7 and 8. What’s the product of 7 and 8? 1) Apply properties of operations as strategies to multiply and divide. $0.7773$
2) Fluently multiply and divide within 100. $0.7549$
Student 56
Tutor Excellent work! Now, don’t forget we still have that $x$ from $7x$. What should we do with it? 1) Write, read, and evaluate expressions in which letters stand for numbers. $0.7773$
Student 56x?
6.3 Learning Curve Analysis
A common method for assessing the quality of KCs and KT methods is examining how well they satisfy pedagogical theory, in particular, the power law of practice [49]: as students practice a particular skill, their mastery of it will increase rapidly at first and then more slowly over time. Following prior work [47], we visualize changes in predicted mastery levels at the KC-level, examine trends in the curves, and investigate whether there is evidence of the power law of practice. We use LLMKT’s predicted masteries on the CoMTA data for all dialogues when they are in the test split of the cross-fold experiment. We use CoMTA instead of MathDial for this analysis since the latter uses LLMs to simulate students; it is unlikely for simulated students to behave like real ones and we confirmed this in a preliminary analysis.
To compute “learning curves”, for a KC $c$, we first find each time that KC occurs in the data, and we use $c^{(i)}_{j}$ to denote its $j^{\text{th}}$ occurrence in the $i^{\text{th}}$ dialogue, and use $\hat{z}^{(i)}_{j}$ to denote the corresponding predicted KC mastery level. We then subtract $\hat{z}^{(i)}_{1}$ from each $\hat{z}^{(i)}_{j}$ to track the change in predicted mastery from the first prediction, since different students may have different initial knowledge levels. We plot the average KC mastery level, i.e., $\frac{1}{N}\sum_{i=1}^{N}\hat{z}^{(i)}_{j}-\hat{z}^{(i)}_{1}$, where $N$ is the number of dialogues that have at least $j$ occurrences of $c$, and show the standard deviation using error bars.
Figure 2: “Learning curves”, i.e., estimated KC mastery levels over dialogue turns, averaged over students for selected KCs in the CoMTA dataset. More common KCs tend to show slightly increasing mastery levels as the dialogue progresses, while less common ones tend to show the opposite. From left to right: “Solve linear equations and inequalities in one variable, including equations with coefficients represented by letters,” “Use variables to represent numbers and write expressions when solving a real-world or mathematical problem,” and “Apply the properties of operations to generate equivalent expressions.”
We examine plots of the 15 most frequent KCs since less common KCs do not have enough data to show clear trends. Out of these, the results are mixed: 7 out of 15 KCs trend towards increasing mastery as the dialogues go on, while 8/15 KCs trend towards decreasing mastery. Figure 2 shows representative examples of each case. KCs with an upward trend typically follow the same pattern: mastery initially has high variance and features a small downward trend, but is then followed by a large and steady increase. Additionally, these KCs are more prevalent in longer dialogues and correspond to the more commonly occurring KCs in the data. These patterns can be explained by two concurrent trends in the data: First, there are several long dialogues where a student struggles with a math topic at the beginning, getting some questions correct and others incorrect, but after some time becomes more proficient and consistently responds correctly; these dialogues contribute to the steady increase in KC mastery. Second, there are several short dialogues where the student either simply knows the topic and responds correctly or struggles with the topic and responds incorrectly before the dialogue quickly ends; these dialogues contribute to the high variance and early dip in KC mastery.
On the contrary, KCs with a downward trend tend to be associated with shorter dialogues and are less frequent in the dataset. These dialogues tend to feature a student repeatedly trying to solve a problem without any success, only to end the dialogue immediately after getting to the correct solution. KC mastery can also drop significantly at the beginning of the dialogue once the student starts to respond incorrectly. Since there are few dialogues to average over, these downward trends can dominate the learning curves; the most common KC with a downward trend only appears in 8 dialogues, which is less than most KCs with upward trends.
Overall, when we take into account the small amount of data, the learning curve trends seem to mostly resemble the power law of practice; when there are sufficient turns in a dialogue, the student’s predicted mastery eventually increases before starting to level out. The main reason these trends are not more prevalent is likely due to the dialogues typically ending after only a few turns. This setting is different than the problem-solving setup studied in existing KT methods, where students practice a KC repeatedly by solving different problems, even after mostly mastering it. Again, we call upon researchers who have access to data with longer dialogues to conduct more extensive analyses.
7 Annotation Quality Experiments
7.1 Human Evaluation Setup and Results
To quantify GPT-4o’s accuracy on the correctness and KC labeling tasks, we conduct an IRB-approved human evaluation using expert human annotators. We hired 3 former math teachers at \$40 US/hour, each rating the same set of 30 randomly selected dialogues from CoMTA with GPT-4o’s labels, totaling 166 turn pairs. For each turn in each dialogue, we ask annotators to score GPT-4o’s label i) on correctness, as 1 if they think the label is accurate and 0 otherwise, and ii) on the standards selected by GPT-4o on a 1-to-4 scale with 4 as the highest. We provide the exact instructions given to annotators in an online Supplementary Material [^4] https://osf.io/wrkej?view_only=f32472a6560649e692a4e2ba603ef52c. For both correctness and KCs, we report the average human-assigned score and two inter-rater reliability (IRR) metrics: i) the portion of times that the labels of all 3 annotators are exactly the same (Overlap) and ii) the Krippendorff’s $\mathbf{\alpha}$ coefficient [5]. We also compute the accuracy of correctness predictions at the final turn using CoMTA’s ground truth labels for both GPT-4o and our annotators.
Table 4 shows the human evaluation results. We see that our annotators give GPT-4o very high scores for correctness and moderate-to-high scores for KCs, indicating that GPT-4o is able to competently perform these tasks. IRR is volatile, with high overlap and low $\alpha$ on correctness, and low overlap but moderate $\alpha$ on KCs. However, the low $\alpha$ on correctness is expected given the high class imbalance of the data (more than 90% of labels are 1), and the low overlap on KCs is expected since there are 4 classes and exact ratings can be subjective. GPT-4o’s accuracy on final turn correctness is lower than the annotator score might suggest, though this is expected since the final turns are much harder to compute correctness on without subsequent tutor turns to help indicate the label. As reference, the CoMTA tech report [35] reports final turn correctness evaluation with GPT-4o at similar accuracy (78.3). Finally, given that the average human annotator’s final turn correctness accuracy is only slightly higher than GPT-4o’s, we can conclude that the correctness labeling task is a challenging one where GPT-4o achieves close to human-level performance.
Table 4: Scores and IRR metrics for the human evaluation. Additionally, final turn correctness accuracy for GPT-4o (on the full dataset followed by only on the human evaluation dataset) and the human annotators.
Label Score Overlap $\alpha$ GPT-4o Final Turn Acc. Human Final Turn Acc.
Correctness 0.9317 / 1 0.8434 0.1809 75.82 (83.33) 85.56
KCs 3.2831 / 4 0.3494 0.4383
7.2 Error Analysis
To understand when GPT-4o’s errors in labeling correctness and KCs happen, we qualitatively analyze its annotations. When labeling correctness, most of GPT-4o’s errors are on the final turn, especially in questions that require some numerical calculations to verify correctness, such as when the student provides the numerical result of a sum or product calculation. This is a known limitation in GPT-4 [45] and can possibly be improved through external tools such as calculators. The other common error tends to be when GPT-4o should label a turn as “na” but does not, e.g., when the tutor asks the student whether they know how to use a function on their calculator. GPT-4o also sometimes labels a turn as correct when the tutor asks a student if they know a concept and the student indicates they do not, showing that GPT-4o can be confused by the tutor’s intentions.
When labeling KCs, GPT-4o’s primary error is not assigning sufficient standards to a turn to describe all the required KCs. This error is mostly due to not selecting enough relevant domains and clusters early in the annotation process, especially for ones with broad or abstract descriptions. As a result, there are not enough related standards to select from to use as KCs. On the other hand, GPT-4o occasionally assigns too many KCs to a turn, selecting all standards that are related to the topic but not necessarily only ones relevant to the current question. These errors can likely be resolved with more sophisticated prompting, potentially running the annotation process multiple times through sampling and taking a majority vote. We leave the exploration of such methods that require more API calls for future work. Finally, there are several dialogues not covered by Common Core standards; there is little that GPT-4o can do other than select a somewhat related standard. This observation suggests that it may be necessary to augment the Common Core standards with other sources to have a more complete set of candidate KCs, or defer to humans for labeling in these cases.
8 Conclusions and Future Work
In this paper, we introduced the task of dialogue KT with the goal of estimating student knowledge from tutor-student dialogues. We used GPT-4o to automatically annotate student response correctness and relevant KCs in the dialogue turns, and showed via an expert human evaluation that these annotations are largely accurate. We introduced a new LLM-based KT method that leverages the textual content in dialogues, and showed that LLMKT i) significantly outperforms existing methods on dialogue KT and ii) works with limited training data. We also showed that our models use the textual content in dialogues to capture student knowledge in meaningful ways.
Considering that this is the first work to attempt KT in dialogues, we acknowledge the limitations of our work but hope it will open up numerous avenues for future work. First, advanced prompting techniques can increase accuracy on the dialogue annotation tasks, and more sophisticated LLM-based methods can improve performance on the KT task. Second, the scope of dialogue KT can be expanded upon: in this work, we only focus on math dialogues between a single student and tutor, but future work can extend to other subjects and dialogues involving multiple students. Third, there are many potential applications of dialogue KT that can be explored. Future work should investigate whether estimated knowledge states are useful for personalization, providing teacher support, or even used to train AI-based tutors by powering simulated student agents. Overall, we hope that this initial investigation can broaden the scope of student modeling to conversational settings, which have become of significant interest lately.
9 Acknowledgement
The authors would like to thank Kristen Dicerbo and the Khan Academy team for guidelines on how to use the CoMTA dataset. We would also like to thank Nigel Fernandez and Zichao Wang for helpful discussions around this work. This work is partially supported by Schmidt Futures under the learning engineering virtual institute (LEVI) initiative and the NSF under grants 2118706, 2237676, and 2341948.